Given:



C= 400+cYd



T=50+.10Y



I=10



G=23



Ye=736.36



The marginal prosperity to consume is

1
Expert's answer
2022-02-03T12:17:25-0500

Y=C+I+G

C=Y-I-G

C=736.36-10-23=703.36

C=400+cYd=400+c(Y-T)

"T=50+0.10Y=50+0.10\times736.36=123.636"

"C=400+cYd=400+c(Y-T)=400+c(736.36-123.636)=400+612.724c"


703.63=400+c612.724

300.36=c612.724

c=0.4951